OpenStudy (lifeisadangerousgame):

This is the question: First, solve the equation for y. (5 points) Second, find the value of y for the given x-value. (5 points) 3x - 2y = -7, when x = 3 This is my answer: Is it correct? 1. 3x - 2y = -7 3x - 2y - 3x = -7 - 3x -2y = -3x + 7 -2y/-2 = -3x + 7/2 y = -3/2x + 7/2 2. 3x - 2y = -7 3 X 3 - 2y = -7 9 - 2y = -7 9 - 2y - 9 = -7 - 9 -2y = -16 y = 8

OpenStudy (anonymous):

Check the second one...why are you multiplying 3 with 3x?

OpenStudy (lifeisadangerousgame):

Because she(my teacher) wants to know what y is and she told me that x is 3

OpenStudy (lifeisadangerousgame):

In the second one, I'm not solving it as a slope equation, but as a regular "find x or y" equation

hero (hero):

Technically, yes, but, I believe it was meant to be solve in a certain way. Check out my solution: 1. 3x - 2y = -7 3x - 2y - 3x = -7 - 3x -2y = -3x - 7 -2y/-2 = 3x + 7/2 y = 3/2x + 7/2 y = (3x + 7)/2 2. y = (3x + 7)/2 y = (3(3)+7)/2 y = (9+7)/2 y = (16)/2 y = 8

hero (hero):

By the way, you also made some algebraic mistakes in the first part. I corrected them for you.

OpenStudy (lifeisadangerousgame):

Ohhh thanks!